OpenStudy (btaylor):

If sin(Θ)=x^2, what is sin(2Θ) in terms of x?

OpenStudy (asnaseer):

do you know how to express sin(2a) in terms of sin(a) and cos(a)?

OpenStudy (btaylor):

um, sin(2a)=2sin(a)cos(a), right?

OpenStudy (asnaseer):

perfect - now you just need to express cos(a) in terms of sin(a). do you know how to do that?

OpenStudy (btaylor):

Is it cos(a)=sin(a+π/2)? cuz that's where i'm stuck.

OpenStudy (asnaseer):

no - use the following identity:\[\sin^2(a)+\cos^2(a)=1\]

OpenStudy (asnaseer):

the rest should be simple

OpenStudy (btaylor):

Oh...OK. So it is 2x^2 sqrt(1-x^4)

OpenStudy (asnaseer):

that looks right to me

OpenStudy (btaylor):

Thanks for your help!

OpenStudy (asnaseer):

yw :)

OpenStudy (zarkon):

note in some cases you will need \[-2x^2\sqrt{1-x^4}\]

OpenStudy (asnaseer):

yes indeed - @Zarkon spotted something we both failed to spot:\[\cos(a)=\pm\sqrt{1-\sin^2(a)}\]